OpenStudy (anonymous):

i need help with this please help 2(40x-10)+45-15=3(25x+15)-5

OpenStudy (anonymous):

2(40x-10)+45-15=3(25x+15)-5 i.e. 80x-40 +30 = 75x + 45 -5 i.e. 80x-10 = 75x + 40 i.e. 80x-75x = 40+10 i.e. 5x = 50 i.e. x= 50/5=10